OpenStudy (anonymous):

Quick series question....Can anyone quickly explain why (k+1)! = (k+1)k!

OpenStudy (jlastino):

n!= n (n-1) (k+1)!=(k+1)(k+1-1)=(k+1)(k) ...wait is there really a factorial sign after k on the right side?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

From what it shows it shows one example before which is K! = k(k-1)....(2)(1)

OpenStudy (anonymous):

and then the next is (k+1)! = (k+1)k!....

OpenStudy (jlastino):

hmm maybe this can help you http://www.themathpage.com/aprecalc/factorial.htm See example 5

OpenStudy (anonymous):

Do you think it would be a fair statement to say that (k+1)k! = (k+1)!k

OpenStudy (callisto):

(k+1)k! = (k+1) [(k)(k-1)(k-2).......x1] = (k+1)! Now what do you think?

OpenStudy (anonymous):

By Definition :- k! = k(k-1)(k-2)....(3)(2)(1) Now substitute k+1 for k in the above definition: (k+1)! = (k+1)(k)(k-1)(k-2)........(3)(2)(1) Observe that after k+1 all the other terms multiply to yield k!. Thus, (k+1)! = (k+1)k!. I hope that answers your query!

OpenStudy (callisto):

But for the statement (k+1)k! = (k+1)!k, I don't think it is correct As I've mentioned (k+1)k! = (k+1) [(k)(k-1)(k-2).......x1] = (k+1)! But for the right hand side, you can see that they are not the same

OpenStudy (anonymous):

Yes, that statement is definitely wrong.

OpenStudy (anonymous):

maybe because it wasn't heard the first time around: (k+1)! nobody hears the shout (k+1) k! somebody prolly heard this one... kidding...:) @Callisto ,@Aron_West are correct....